# Beyond the Doors of Perception: Vision Transformers Represent Relations Between Objects

### v1

@@ 5 The Relational Stage
We now characterize the relational stage, where attention between tokens largely excludes other tokens within the same object. We hypothesize that this stage takes in the object representations formed by the perceptual stage and computes relational same-different operations over them. We find that the operations implemented by these relational layers are somewhat abstract in that 1) they do not rely on memorizing individual objects and 2) one can identify abstract same and abstract different representations in the RMTS task, which are constant even as the perceptual qualities of the underlying object pairs vary.
@@
Figure 5: Linear probing and intervention results. We probe for the intermediate same-different judgments required to perform the RMTS task (blue). We see that probe performance reaches ceiling at layer 5 and maintains throughout the rest of the network. We use the directions defined by the linear probe to intervene on model representations in order to flip an intermediate judgment (green). We see that this intervention succeeds reliably at layer 5 and then drops. We add a vector that is consistent with a pair’s exhibited same-different relation as a control, and see that this has little effect (yellow). The background is colored according to the heatmap in Figure 2(c).
@@
Methods – Patching Novel Representations: In Section 4, we identify independent linear subspaces encoding shape and color information. We now ask: does the content contained in these subspaces matter to the overall same-different computation? One can imagine an ideal version of the same-different relation that is abstracted away from the particular properties of the objects it is comparing. In this setting, the same-different relation could take in object representations where colors and shapes are represented by any arbitrary vector, and the model still correctly judges sameness or difference. To study this, we intervene on either the linear subspaces for shape or color for both objects in a pair, and replace the content found therein with novel representations (See Figure 4(a)). To create a “different” example, we start with a “same” image and replace the shape (or color) representations of both objects with two different novel representations, and to create a “same” example, we start with a “different” image and replace them with two identical novel representations. We then assess whether the model’s decision changes accordingly. We generate novel representations using four methods: 1) we add the representations found within the linear subspaces corresponding to two randomly sampled objects in an IID Validation set, 2) we interpolate between these representations, 3) we sample each dimension randomly from a distribution of embeddings, and 4) we sample each dimension from an OOD random distribution (a mean 0 normal distribution). See Appendix H for technical details.
@@
Results: The results of patching in novel representations are presented in Figure 4(b). Overall, we find the greatest success when patching in added representations, followed by interpolated representations. We observe limited success when patching sampled representations, and no success patching random mean 0 vectors. All interventions perform best in layers 2 and 3, towards the end of the perceptual stage. Overall, this points to a limited form of abstraction present in the relational stage — the same-different operation is not fully abstract (in that it cannot operate over arbitrary vectors), but it does not rely on rote memorization of specific object representations (as it can generalize to additions and interpolations of these representations).
@@
Methods – Linear Interventions: The RMTS task allows us to further characterize the relational stage, as it requires forming an intermediate representation of same and different, and then comparing them. In this analysis, we ask: are these intermediate representations invariant to the perceptual qualities of the object pairs that underlie them? We linearly probe for intermediate same or different judgments from the collection of tokens corresponding to object pairs. The probe consists of a linear transformation mapping the residual stream to 2 dimensions, representing same and different. Each row of this transformation can be viewed as a direction $d$ in the residual stream corresponding to the value that you are probing for (i.e. $d_{same}$ is the linear direction representing $same$). We train one probe for each layer on images from the model’s train set, and test on images from its test set. To understand whether the directions discovered by the probe are causally implicated in model behavior, we (once again) create a counterfactual intervention Nanda et al. (2023). In order to change an intermediate judgment from same to different, we merely add the direction $d_{diff}$ to the intermediate representations of objects that exhibit the same relation, and observe whether the model behaves as if this pair now exhibits the different relation [^6] Prior work normalizes the intervention directions and searches over a scaling parameter. In our setting, we find that simply adding the direction defined by probe weights without rescaling works well, though see Appendix I for further exploration. We use transformerlens for this intervention Nanda & Bloom (2022).. We run this intervention on images from the model’s test set. We also run a control intervention, where we add the incorrect direction (i.e., we add the $d_{same}$ when the object pair already expressed the “same” relation). We anticipate that this will not reliably flip the model’s downstream decisions.
@@
Results: Results of probing and linear interventions are shown in Figure 5. We observe that linear probe performance peaks in the middle layers of the model (layer 5) and then stays high. However, our linear intervention counterfactual accuracy peaks at layer 5 and then drops precipitously. Notably, layer 5 also corresponds to the peak of within-pair attention (See Figure 2. This indicates that, at least in layer 5, there exists a single direction representing same and a single direction representing different. One can flip the intermediate same-different judgment by adding a vector in one of these directions to the residual streams of any pair of objects. Finally, the control intervention completely fails throughout all layers, as expected. Thus, the CLIP-pretrained ViT does in fact generate abstract representations of same and different, and operate on them in order to produce an RMTS response.
@@
Figure 6: We average the best counterfactual intervention accuracy for shape and color and plot it against IID, OOD, and Compositional Test set performance for CLIP, DINO, ImageNet, and from-scratch B/16 models. We observe that increased disentanglement (i.e. higher counterfactual accuracy) correlates with downstream performance. The from-scratch model achieved only chance IID performance in RMTS, so we omitted it from the analysis.
@@ 6 Disentanglement Correlates with Generalization Performance
Having object representations that disentangle perceptual properties may enable a model to generalize to out-of-distribution stimuli. As prior work has argued, having disentangled visual representations should enable compositional generalization to unseen combinations of perceptual properties (Higgins et al. (2018); Bengio et al. (2013), cf. Locatello et al. (2019)). To investigate the relationship between disentanglement and generalization, we finetune CLIP, ImageNet, DINO, and randomly-initialized models on a new dataset, where each shape is only ever paired with two distinct colors. We then repeat our analyses in Section 4 to identify independent linear subspaces for shape and color [^7] In order to train these interventions, we must generate counterfactual images. These counterfactual images contain shape-color pairs that are not seen during model training. However, we emphasize that we only use these interventions to derive a disentanglement metric (which we expect to be correlated with compositional generalization), not to train model weights.. We evaluate models in 3 settings: on an IID test set consisting of observed shape-color combinations, on a compositional generalization test set consisting of unobserved shape-color combinations (where each shape and each color have been individually observed), and finally on an OOD consisting of completely novel shapes paired with completely novel colors. We plot the relationship between disentanglement (i.e. counterfactual intervention accuracy) and overall performance on these model evaluations. We find a consistent trend: more disentangled representations correlated with downstream model performance in all cases (See Figure 6).
@@ 7 Failure Modes
Previous sections have argued that ViTs trained on same-different tasks implement a two-stage processing pipeline to arrive at a response. In this section, we argue that both perceptual and relational stages could serve as failure points for models, impeding their ability to solve a same-different task. In practice, tasks that rely on relations between objects likely have perceptual and relational stages that are orders of magnitude more complex than those we study here. The results presented herein indicate that solutions targeting either the perceptual (Zeng et al., 2022) or relational (Bugliarello et al., 2023) stages may be insufficient for producing the robust, abstract computations that we desire.
Perceptual Failure Modes: First, we note that models trained from scratch on a discrimination task do not clearly distinguish between perceptual and relational stages (See Figure 2(b)). As Table 1 shows, a model trained on a limited number of shape-color combinations model achieves high train accuracy but poor test accuracy. We hypothesize that a failure to produce object-level representations is precluding the model from solving the discrimination task. To test, we introduce an auxiliary loss to encourage models to form disentangled object-level representations. During training, we introduce token-level probes that are optimized to predict shape information from one linear subspace (i.e., the first 384 dimensions) of the representations generated at an intermediate layer of the model and color information from the complementary linear subspace at that same layer. These probes are optimized during training, and the probing loss is backpropagated through the model. From Table 1, we see that this auxiliary loss enables high test accuracy and also improve compositional generalization.
Aux. Loss Task Patch Size Train Acc. Test Acc. Comp. Acc.
Discrimination 16 77.3% 76.5% 75.9%
Discrimination 16 97.3% 94.6% 86.1%
Discrimination 32 99.1% 88.0% 78.3%
Discrimination 32 99.7% 99.7% 93.5%
@@
Table 1: Adding an auxiliary loss term to encourage disentangled object representations boosts test accuracy and compositional generalization in models trained from scratch on the discrimination task.
Relational Failure Modes: We might ask whether the same approach can help models trained from-scratch on the RMTS task – models that achieve only chance performance. Despite our auxiliary loss probes achieving high IID probing accuracy, our intervention still results in chance test accuracy (See Appendix J). This provides evidence that 1) models may fail at either the perceptual or relational stages and 2) that the relation defined by the discrimination task is simple to solve with disentangled object representations, while the relation defined in the RMTS task is still far from trivial.
@@ 8 Discussion
Related Work: This work takes inspiration from the field of mechanistic interpretability, which seeks to characterize the algorithms that neural networks implement (Olah, 2022). Though many of these ideas originated in the domain of NLP (Wang et al., 2022; Hanna et al., 2024; Feng & Steinhardt, 2023; Wu et al., 2024c; Merullo et al., 2023; Geva et al., 2022; Meng et al., 2022) and in toy settings (Nanda et al., 2022; Elhage et al., 2022; Li et al., 2022), they are beginning to find applications in computer vision (Fel et al., 2023; Vilas et al., 2024; Palit et al., 2023). These techniques augment an already-robust suite of tools that visualize the features (rather than algorithms) that vision models use (Olah et al., 2017; Selvaraju et al., 2017; Simonyan et al., 2014). Finally, this study contributes to a growing literature employing mechanistic interpretability to address debates within cognitive science (Millière & Buckner, 2024; Lepori et al., 2023a; Kallini et al., 2024; Traylor et al., 2024).
@@
Conclusion: The ability to compute abstract visual relations is a fundamental aspect of biological visual intelligence and a crucial stepping stone toward useful and robust artificial vision systems. In this work, we demonstrate that some finetuned ViTs adopt a two-stage processing pipeline to solve same-different tasks – despite having no obvious inductive biases towards this algorithm. First, models produce disentangled object representations in a perceptual stage; then models compute a somewhat abstract version of the same-different computation in a relational stage. Finally, we observed a correlation between disentanglement and generalization and noted that models might fail to learn either the perceptual or relational operations necessary to solve a task. One question raised by this study is why CLIP-pretrained models perform favorably (and adopt this two-algorithm so cleanly) relative to other pretrained models. Raghu et al. (2021) finds that models pretrained on more data tend to learn local attention patterns in early layers, followed by global patterns in later layers. This might give CLIP a particular incentive to create local object representations, which are then used in relational operations. Future work might test this hypothesis. Future work might also focus on pinning down the precise mechanisms (i.e., the attention heads and MLPs) used to implement the perceptual and relational stages or on generalizing these findings to more complex datasets.
Appendix A Constructing the Objects
Figure A demonstrates a single instance of the $16$ shapes and $16$ colors used in our datasets. Any shape can be paired with any color to create one of $256$ unique objects. Note that object colors are not uniform within a given object. Instead, each color is defined by three different Gaussian distributions—one for each RGB channel in the image—that determine the value of each object pixel. For example, the color red is created by these three distributions: $\mathcal{N}(\mu=233,\sigma=10)$ in the red channel, $\mathcal{N}(\mu=30,\sigma=10)$ in the green channel, and $\mathcal{N}(\mu=90,\sigma=10)$ in the blue channel. All color distributions have a variance fixed at $10$ to give them an equal degree of noise. Any sampled values that lie outside of the valid RGB range of $[0,255]$ are clipped to either 00 or $255$. Object colors are re-randomized for every image, so no two objects have the same pixel values even if they are the same color. This was done to prevent the models from learning simple heuristics like comparing single pixels in each object.
@@ Appendix A Constructing the Objects
Figure 7: All 16 unique shapes and colors used to construct the Discrimination and RMTS tasks.
Appendix B ViT B/16 All Model Results

### v3

@@ 5 The Relational Stage
We now characterize the relational stage, where tokens within one object largely attend to tokens in the other object. We hypothesize that this stage takes in the object representations formed in the perceptual stage and computes relational same-different operations over them. We find that the operations implemented by these relational layers are somewhat abstract in that 1) they do not rely on memorizing individual objects and 2) one can identify abstract same and different representations in the RMTS task, which are constant even as the perceptual qualities of the object pairs vary.
@@
Figure 5: Linear probing and intervention results. We probe for the intermediate same-different judgments required to perform the RMTS task (blue). Probe performance reaches ceiling at around layer 5 and maintains throughout the rest of the model. We use the directions defined by the linear probe to intervene on model representations and flip an intermediate judgment (green). This intervention succeeds reliably at layer 5 but not deeper. We add a vector that is consistent with a pair’s exhibited same-different relation as a control (yellow). This has little effect. The background is colored according to the heatmap in Figure 2c (blue$=$local heads; red$=$global heads).
@@
Methods — Patching Novel Representations. In Section 4, we identify independent linear subspaces encoding shape and color. Does the content of these subspaces matter to the same-different computation? One can imagine an ideal same-different relation that is completely abstracted away from the particular properties of the objects being compared. In this setting, a model could accurately judge “same” vs. “different” for object representations where colors and shapes are represented by arbitrary vectors. To study this, we intervene on the linear subspaces for either shape or color for both objects in a pair, replacing the content found therein with novel representations (see Figure 4a). To create a “different” example, we start with a “same” image and replace the shape (or color) representations of both objects with two different novel representations; to create a “same” example, we start with a “different” image and replace them with two identical novel representations. We then assess whether the model’s decision changes accordingly. We generate novel representations using four methods: 1) we add the representations found within the linear subspaces corresponding to two randomly sampled objects in an IID validation set, 2) we interpolate between these representations, 3) we sample each dimension randomly from a distribution of embeddings, and 4) we sample each dimension from an OOD random distribution (a $\mu=0$ normal). See Appendix J for technical details.
@@
Results. The results of patching novel representations into a CLIP-pretrained ViT are presented in Figure 4b. Overall, we find the greatest success when patching in added representations, followed by interpolated representations. We observe limited success when patching sampled representations, and no success patching random vectors. All interventions perform best in layers 2 and 3, towards the end of the perceptual stage. Overall, this points to a limited form of abstraction in the relational stage. The same-different operation is somewhat abstract—while it cannot operate over completely arbitrary vectors, it can generalize to additions and interpolations of shape & color representations, indicating that it does not rely on rote memorization of specific objects. Results for CLIP-pretrained ViTs on RMTS and other models on both tasks are found in Appendix K. Results for DINOv2 are found in Appendix D. Other models largely produce similar results to CLIP in this analysis.
@@
Methods — Linear Interventions. The RMTS task allows us to further characterize the relational stage, as it requires first forming then comparing intermediate representations of same and different. Are these intermediate representations abstract (i.e. invariant to the perceptual qualities of the object pairs that underlie them)? We linearly probe for intermediate same or different judgments from the collection of tokens corresponding to object pairs. The probe consists of a linear transformation mapping the residual stream to two dimensions representing same and different. Each row of this transformation can be viewed as a direction $d$ in the residual stream corresponding to the value being probed for (e.g. $d_{\text{same}}$ is the linear direction representing same). We train one probe for each layer on images from the model’s train set and test on images from a test set. To understand whether the directions discovered by the probe are causally implicated in model behavior, we create a counterfactual intervention (Nanda et al., 2023). In order to change an intermediate judgment from same to different, we add the direction $d_{\text{diff}}$ to the intermediate representations of objects that exhibit the same relation. We then observe whether the model behaves as if this pair now exhibits the different relation. [^6] Prior work normalizes the intervention directions and searches over a scaling parameter. In our setting, we find that simply adding the direction defined by probe weights without rescaling works well. See Appendix K for further exploration. We use transformerlens for this intervention (Nanda & Bloom, 2022).
We run this intervention on images from the model’s test set. We also run a control intervention where we add the incorrect direction (e.g., we add $d_{\text{same}}$ when the object pair is already “same”). This control intervention should not reliably flip the model’s downstream decisions.
@@
Results. Probing and linear intervention results for a CLIP-pretrained ViT are shown in Figure 5. We observe that linear probe performance peaks in the middle layers of the model (layer 5) and then remains high. However, our linear intervention accuracy peaks at layer 5 and then drops precipitously. Notably, layer 5 also corresponds to the peak of within-pair attention (see Figure 2c). This indicates that—at least in layer 5—there exists a single direction representing same and a single direction representing different. One can flip the intermediate same-different judgment by adding a vector in one of these directions to the residual streams of any pair of objects.
Finally, the control intervention completely fails throughout all layers, as expected. Thus, CLIP ViT does in fact generate and operate over abstract representations of same and different in the RMTS task. We find similar results for a DINOv2 pretrained model (see Appendix D), but not for others (see Appendix K).
@@
Figure 6: We average the best counterfactual intervention accuracy for shape and color and plot it against IID, OOD, and Compositional Test set performance for CLIP, DINO, DINOv2, ImageNet, MAE, and from-scratch B/16 models. We observe that increased disentanglement (i.e. higher counterfactual accuracy) correlates with downstream performance. The from-scratch model achieved only chance IID performance in RMTS, so we omitted it from the analysis.
@@ 6 Disentanglement Correlates with Generalization Performance
Object representations that disentangle perceptual properties may enable a model to generalize to out-of-distribution stimuli. Specifically, disentangled visual representations may enable compositional generalization to unseen combinations of perceptual properties (Higgins et al. (2018); Bengio et al. (2013), cf. Locatello et al. (2019)) [^7] Though Locatello et al. (2019) find that then-current measurements of of disentanglement fail to correlate with downstream performance in variational autoencoders, many of those models failed to produce disentangled representations in the first place.. To investigate the relationship between disentanglement and generalization, we fine-tune CLIP, ImageNet, DINO, DINOv2, MAE, and randomly-initialized ViTs on a new dataset where each shape is only ever paired with two distinct colors. We then repeat our analyses in Section 4 to identify independent linear subspaces for shape and color. [^8] In order to train these interventions, we generate counterfactual images, some of which contain shape-color pairs that are not seen during model training. However, we emphasize that we only use these interventions to derive a disentanglement metric, not to train model weights. We evaluate models in 3 settings: 1) on an IID test set consisting of observed shape-color combinations, 2) on a compositional generalization test set consisting of unobserved shape-color combinations (where each shape and each color have been individually observed), and 3) on an OOD test set consisting of completely novel shapes and colors. We plot the relationship between disentanglement (i.e. counterfactual intervention accuracy) and overall performance on these model evaluations. We find a consistent trend: more disentangled representations correlates with downstream model performance in all cases (See Figure 6).
@@ 7 Failure Modes
Previous sections have argued that pretrained ViTs that achieve high performance when finetuned on same-different tasks implement a two-stage processing pipeline. In this section, we argue that both perceptual and relational stages can serve as failure points for models, impeding their ability to solve same-different tasks. In practice, tasks that rely on relations between objects likely have perceptual and relational stages that are orders of magnitude more complex than those we study here. The results presented herein indicate that solutions targeting either the perceptual (Zeng et al., 2022) or relational (Bugliarello et al., 2023) stages may be insufficient for producing the robust, abstract computations that we desire.
Perceptual and Relational Regularizers. We introduce two loss functions, designed to induce disentangled object representations and multi-stage relational processing, respectively. When employing the disentanglement loss, we introduce token-level probes that are optimized to predict shape information from one linear subspace (e.g., the first 384 dimensions) of the representations generated at an intermediate layer of the model and color information from the complementary linear subspace at that same layer (layer 3, in our experiments). These probes are optimized during training, and the probe loss is backpropagated through the model. This approach is motivated by classic work on disentangled representations (Eastwood & Williams, 2018). The pipeline loss is designed to encourage discrete, specific stages of processing by regularizing the attention maps to maximize the attention pattern scores defined in Section 3. Specifically, early layers are encouraged to maximize attention within-object, then within-pair, and finally (in the case of RMTS stimuli) between-pair. See Appendix L for technical details. Note that the disentanglement loss targets the perceptual stage of processing, whereas the pipeline loss targets both perceptual and relational stages.
Results. First, we note that models trained from scratch on the discrimination task do not clearly distinguish between perceptual and relational stages (Figure 2b). Thus, we might expect that a model trained on a limited number of shape-color combinations would not learn a robust representation of the same-different relation. Indeed, Table 1 confirms this. However, we see that either the disentanglement loss or the pipeline loss is sufficient for learning a generalizable representation of this relation.
Similarly, we find that models trained from scratch on the RMTS task only achieve chance performance. However, in this case we must include both disentanglement and pipeline losses in order to induce a fairly general (though still far from perfect) hierarchical representation of same-different. This provides evidence that models may fail at either the perceptual or relational stages: they might fail to produce the correct types of object representations, and/or they might fail to execute relational operations over them. See Appendix 7 for further analysis.
Task Disent. Loss Pipeline Loss Train Acc. Test Acc. Comp. Acc.
Disc. 77.3% 76.5% 76.0%
Disc. 97.3% (+20) 94.6% (+18.1) 86.1% (+10.1)
Disc. 95.6% (+18.3) 93.9% (+17.4) 92.3% (+16.4)
RMTS 49.2% 50.1% 50.0%
RMTS 53.9% (+4.7) 54.4% (+4.3) 54.1% (+4.1)
RMTS 66.1% (+16.9) 50.1% 50.1% (+0.1)
RMTS 95.1% (+45.9) 94.1% (+44) 77.4% (+27.4)
@@
Table 1: Performance of ViTs trained from scratch with auxiliary losses. Adding either a disentanglement loss term to encourage disentangled object representations (Disent. Loss) or a pipeline loss to encourage two-stage processing in the attention heads (Pipeline Loss) boosts test accuracy and compositional generalization (Comp. Acc.) for the discrimination task. Both auxiliary losses are required to boost accuracy for the RMTS task.
@@ 8 Discussion
Related Work. This work takes inspiration from the field of mechanistic interpretability, which seeks to characterize the algorithms that neural networks implement (Olah, 2022). Though many of these ideas originated in the domain of NLP (Wang et al., 2022; Hanna et al., 2024; Feng & Steinhardt, 2023; Wu et al., 2024c; Merullo et al., 2023; Geva et al., 2022; Meng et al., 2022) and in toy settings (Nanda et al., 2022; Elhage et al., 2022; Li et al., 2022), they are beginning to find applications in computer vision (Fel et al., 2023; Vilas et al., 2024; Palit et al., 2023). These techniques augment an already-robust suite of tools that visualize the features (rather than algorithms) that vision models use (Olah et al., 2017; Selvaraju et al., 2017; Simonyan et al., 2014). Finally, this study contributes to a growing literature employing mechanistic interpretability to address debates within cognitive science (Millière & Buckner, 2024; Lepori et al., 2023a; Kallini et al., 2024; Traylor et al., 2024).
@@
Conclusion. The ability to compute abstract visual relations is a fundamental aspect of biological visual intelligence and a crucial stepping stone toward useful and robust artificial vision systems. In this work, we demonstrate that some fine-tuned ViTs adopt a two-stage processing pipeline to solve same-different tasks—despite having no obvious inductive biases towards this algorithm. First, models produce disentangled object representations in a perceptual stage; models then compute a somewhat abstract version of the same-different computation in a relational stage. Finally, we observe a correlation between disentanglement and generalization and note that models might fail to learn either the perceptual or relational operations necessary to solve a task.
Why do CLIP and DINOv2-pretrained ViTs perform favorably and adopt this two-stage algorithm so cleanly relative to other pretrained models? Raghu et al. (2021) find that models pretrained on more data tend to learn local attention patterns in early layers, followed by global patterns in later layers. Thus, pretraining scale (rather than training objective) might enable these models to first form local object representations, which are then used in global relational operations. Future work might focus on pinning down the precise relationship between data scale and relational reasoning ability, potentially by studying the training dynamics of these models. Additionally, future work might focus on characterizing the precise mechanisms (e.g. the attention heads and MLPs) used to implement the perceptual and relational stages, or generalize our findings to more complex relational tasks.
Appendix A Dataset Details
@@
Figure 7: All 16 unique shapes and colors used to construct the Discrimination and RMTS tasks. There are thus $16\times 16=256$ unique objects in our same-different datasets.
A.1 Constructing the Objects
Figure 7 demonstrates a single instance of the $16$ shapes and $16$ colors used in our datasets. Any shape can be paired with any color to create one of $256$ unique objects. Note that object colors are not uniform within a given object. Instead, each color is defined by three different Gaussian distributions—one for each RGB channel in the image—that determine the value of each object pixel. For example, the color red is created by these three distributions: $\mathcal{N}(\mu=233,\sigma=10)$ in the red channel, $\mathcal{N}(\mu=30,\sigma=10)$ in the green channel, and $\mathcal{N}(\mu=90,\sigma=10)$ in the blue channel. All color distributions have a variance fixed at $10$ to give them an equal degree of noise. Any sampled values that lie outside of the valid RGB range of $[0,255]$ are clipped to either 00 or $255$. Object colors are re-randomized for every image, so no two objects have the same pixel values even if they are the same color. This was done to prevent the models from learning simple heuristics like comparing single pixels in each object.
Figure 8: More examples of stimuli for the discrimination and RMTS tasks. The top row shows “different” examples, while the bottom row shows “same” examples. Note that “different” pairs may differ in one or both dimensions (shape & color).
A.2 Constructing the Datasets
The train, validation, and test sets for both the discrimination and RMTS tasks each contain $6,400$ unique stimuli: $3,200$ “same” and $3,200$ “different.” To construct a given dataset, we first generate all possible same and different pairs of the $256$ unique objects (see Figure 7). We consider two objects to be the same if they match in both shape and color—otherwise, they are different. Next, we randomly select a subset of the possible object pairs to create the stimuli such that each unique object is in at least one pair. For the RMTS dataset, we repeat this process to select same and different pairs of pairs.
Each object is resized (from $224\times 244$ pixel masks of the object’s shape) such that it is contained within a single ViT patch for B/32 models or four ViT patches for B/16 & B/14 models. For B/32 and B/16 models, objects are roughly $28\times 28$ pixels in size; for B/14 models (DINOv2 only), objects are roughly $21\times 21$ pixels in size. These choices in size mean that a single object can be placed in the center of a $32\times 32$ (or $28\times 28$) pixel patch with a radius of $4$ pixels of extra space around it. This extra space allows us to randomly jitter object positions within the ViT patches.
To create a stimulus, a pair of objects is placed over a $224\times 224$ pixel white background in randomly selected, non-overlapping positions such that objects are aligned with ViT patches. For RMTS stimuli, the second “display” pair is always placed in the top left corner of the image. Each object’s position (including the “display” objects for RMTS) is also randomly jittered within the ViT patches it occupies. We consider two objects in a specific placement as one unique stimulus—in other words, a given pair of objects may appear in multiple images but in different positions. All object pairs appear the same number of times to ensure that each unique object is equally represented.
See Figure 8 for some more examples of stimuli from each task.
Figure 9: Examples of stimuli from our photorealistic same-different evaluation dataset. The top row contains “different” examples, while the bottom row contains “same” examples. Stimuli are constructed using 16 unique 3D models of objects placed on a table with a randomized texture; background textures are also randomized. Objects are randomly rotated and may be placed at different distances from the camera or occlude each other.
A.3 Photorealistic Test Set
In order to ensure the robustness of the two-stage processing we observe in CLIP and DINOv2 on our artificial stimuli, we test models on a highly out-of-distribution photorealistic discrimination task. The test dataset consists of $1,024$ photorealistic same-different stimuli that we generated (see Figure 9). Each stimulus is a 224$\times$224 pixel image depicting a pair of same or different 3D objects arranged on the surface of a table in a sunlit room. We created these images in Blender, a sophisticated 3D modeling tool, using a set of 16 unique 3D models of different objects that vary in shape, texture and color. To construct the dataset, we first generate all possible pairs of same or different objects, then select a subset of the possible “different” pairs such that each object appears in two pairs. This ensures that all objects are equally represented and that an equal number of “same” and “different” stimuli are created. We create 32 unique stimuli for each pair of objects by placing them on the table in eight random configurations within the view of four different camera angles, allowing partial occlusions. Each individual object is also randomly rotated around its $z$-axis in each image—because 11 of the objects lack rotational symmetry, these rotations provide an additional challenge, especially for “same” classifications.
Figure 10: Attention pattern analysis for CLIP and DINOv2 on the photorealistic discrimination task. This figure follows the top row in Figure 2. (a) CLIP: As in Figure 2, WO peaks at layer 3, WP peaks at layer 6, and BG peaks at layer 10. BG attention is higher throughout the perceptual stage, leading to a lower perceptual score compared to the artificial discrimination task (i.e. fewer blue cells). (b) DINOv2: The attention pattern exhibits two stages, resembling the artificial setting (although the correspondence is somewhat looser than CLIP’s, perhaps explaining DINOv2’s poor zero-shot performance on the photorealistic task).
We evaluate models that have been fine-tuned on the discrimination task from the main body of the paper (e.g. Figure 8a) in a zero-shot manner on the photorealistic dataset, meaning that there is no additional fine-tuning on the photorealistic dataset. We find that CLIP ViT attains a test accuracy of $93.9$% on the photorealistic dataset, while all other models attain chance level accuracy (e.g. DINOv2 attains an accuracy of $48$%). We also find that CLIP performs two-stage processing on the photorealistic stimuli (see Figure 10a), and that the peaks in WO, WP, and BG attention all occur at the same exact layers as the artificial stimuli (i.e. in Figure 2). DINOv2 also displays similar two-stage processing despite its poor performance on the photorealistic task (see Figure 10b). Note that BG attention for both models is higher overall during the perceptual stage when processing the photorealistic stimuli compared to the artificial stimuli; this is likely because the photorealistic stimuli contain detailed backgrounds, while the backgrounds in the artificial stimuli are blank. Overall, these findings generalize our results from the toy setting presented in the main body of the paper.
Appendix B ViT B/16: All Model Behavioral Results
See Tables 2,  3,  4, and 5 for behavioral results from all ViT-B/16 models trained on discrimination and RMTS tasks with either all 256 shape-color combinations or only 32 shape-color combinations. The “Pretraining Scale” column denotes the number of images (in millions) in a given model’s pretraining dataset. The models are organized in descending order by pretraining scale. “Test Acc.” refers to IID test accuracy. “Comp. Acc.” refers to compositional generalization accuracy (for models trained on only 32 shape-color combinations). “Realistic Acc.” (Table 2 only) refers to a model’s zero-shot accuracy on the photorealistic evaluation dataset. CLIP and DINOv2—the two models with a pretraining scale on the order of 100 million images—attain near perfect test accuracy on the RMTS task. However, only CLIP attains high performance on the photorealistic dataset.
Pretrain Pretraining Scale $\downarrow$ Train Acc. Test Acc. Realistic Acc.
CLIP 400M 100% 99.3% 93.9%
DINOv2 142M 100% 99.5% 48.0%
ImageNet 14.2M 100% 97.5% 53.0%
DINO 1.28M 100% 95.6% 50.9%
MAE 1.28M 100% 98.0% 52.4%
95.9% 80.5% 49.9%